# Compatibility of canonical $\ell$-adic local systems on Shimura varieties

### v1

@@ 1. Introduction
    For all closed points $x\in\mathcal{S}_{K_{0},s}[1/\ell]$, the class of $\rho_{\ell}(\mathrm{Frob}_{x})$ in $[G{\,\!\sslash\!\,}G](\overline{\mathbb{Q}}_{\ell})$ lies in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$ and is independent of $\ell$ (not equal to the residue characteristic of $x$). [^1] In fact, $\rho_{\ell}(\mathrm{Frob}_{x})$ is semisimple, by [3, Theorem 7.1]
@@
Theorem 1.1 results from combining our previous work [16] with the work of Bakker-Shankar-Tsimerman ([3]) showing that all Shimura varieties have canonical integral models away from a finite set of bad primes (essentially the $N$ of Theorem 1.1). We should clarify the logical relation between these three papers: [3] makes use of Part (1) of Theorem 1.1 (see Lemma 5.1, Theorem 5.2 of loc. cit.), which we established for $\rho_{\ell}^{\mathrm{ad}}$ in [16, Theorem 2.3]. In §3, we give a proof of Part (1) of Theorem 1.1 in general that does not depend on the two earlier papers. Using their main result on integral models along with our main theorem from [16] (the analogue of Part (2) of Theorem 1.1 for $\rho_{\ell}^{\mathrm{ad}}$), Bakker-Shankar Tsimerman prove ([3, Theorem 9.7]) that for $v\nmid N$, every $\mu$-ordinary point of the special fiber $\mathcal{S}_{v}$ lifts to a characteristic zero special point. [^2] To be precise, [3] use our earlier result for $\rho_{\ell}^{\mathrm{ad}}$ to cover the non-abelian type (or the real rank at most 1) cases; in abelian type, [13, Corollary 2.3.1] establishes the desired compatibility. We in turn use this result on CM lifting of $\mu$-ordinary points to establish Parts (2) and (3) of Theorem 1.1.
@@
We now summarize each section and the arguments therein. In §2, we recall results of Drinfeld ([7]) on Deligne’s companions conjecture; we need his work on companions for representations valued in not necessarily connected reductive (not only semisimple) groups, and so we elaborate slightly on the results in [7, §6]. In §3, we show (Corollary 3.4) that Part (1) of Theorem 1.1 holds: there is an integer $N$ such that for all $\ell$, all $\rho_{\ell}$ are unramified away from $N\ell$. For the $G^{\mathrm{ad}}$-representations, we showed this in [16, Proposition 2.3] by an abstract argument using superrigidity and Deligne’s conjecture. When $G$ is not an adjoint group, descents to $\pi_{1}(S_{K_{0},s})$ of a given representation of $\pi_{1}^{\mathrm{top}}(S_{K_{0},s}(\mathbb{C}))$ (extended to $\pi_{1}((S_{K_{0},s})_{\overline{\mathbb{Q}}})$ using integrality) are not unique, and it is easy to choose descents for varying $\ell$ that have no common finite bad set $N$ as in Part (1) of the Theorem. We need to use the fact that our descents arise from the family of canonical models; it turns out that all we need in addition to the arguments in the adjoint case is that there is a number field point $y$ of $S_{K_{0},s}$ for which the specialized local systems $\rho_{\ell}|_{y}$ are unramified outside $N_{y}\ell$ for some integer $N_{y}$ independent of $\ell$. In §4 we prove Parts (2) and (3) of Theorem 1.1. Given the result of §3, from [3, Theorem 9.2] we deduce that on the $\mu$-ordinary locus of the special fiber $\mathcal{S}_{v}$ ($v\nmid N\ell\ell^{\prime}$), $\rho_{\ell}$ and $\rho_{\ell^{\prime}}$ satisfy the companions property (2). Using the companions construction (§2) and Zariski-density of the $\mu$-ordinary locus, we can deduce the companions property on all of $\mathcal{S}_{v}$ via a strong multiplicity 1 property for Galois representations. The crystalline compatibility works analogously but requires as additional input the overconvergent ($G$-valued) F-isocrystals associated to the crystalline (by Esnault-Groechenig: [19, Appendix]) local systems $\rho_{p}$ in our joint work [9].
@@
We end this introduction by remarking on what we do not do, and what we hope to pursue in future work. We view our results in the framework of the Langlands-Rapoport conjecture, as a significant step toward associating Kottwitz triples to mod $p$ points on Shimura varieties; in abelian type, Kisin has proven the Langlands-Rapoport conjecture ([13]), but this remains a wide-open and fundamental problem in non-abelian type. We further note that Kottwitz triples are not only a useful intermediary in work on the full Langlands-Rapoport conjecture (as in loc. cit.) but are in principle a sufficient intermediary to calculate the zeta function of the Shimura variety; indeed, this is how Kottwitz first employed them in [14]. Recall ([13, §4.3]) that a Kottwitz triple over $k=\mathbb{F}_{p^{r}}$ is a triple $(\gamma_{0},(\gamma_{\ell})_{\ell\neq p},\delta)$ where
@@
  • After possibly replacing $k$ by a finite extension, there is an inner twisting $I$ of $I_{0}:=\mathrm{Cent}_{G}(\gamma_{0})$ such that $I\otimes_{\mathbb{Q}}\mathbb{R}$ is anisotropic mod center, and for all primes $\ell$, $I\otimes_{\mathbb{Q}}\mathbb{Q}_{\ell}$ is isomorphic to $I_{\ell}$ as an inner twist of $I_{0}\otimes_{\mathbb{Q}}\mathbb{Q}_{\ell}$ (here $I_{\ell}$ for $\ell\neq p$ is $\mathrm{Cent}_{G_{\mathbb{Q}_{\ell}}}(\gamma_{\ell})$, and for $\ell=p$ it is the Frobenius-centralizer: see [13, §4.3.1] for details).
To every $k$-point $x\in\mathcal{S}(k)$ ($p\nmid N$) we associate a triple $(\gamma_{0},(\gamma_{\ell})_{\ell\neq p},\delta)$, except that we have only constructed $\gamma_{0}$ in general as an element of $[G{\,\!\sslash\!\,}G](\mathbb{Q})$. With that caveat, we have verified the first and second desiderata for the triple: $\gamma_{\ell}$ is of course $\rho_{\ell}(\mathrm{Frob}_{x})$, and $\delta$ arises from the absolute Frobenius ($\gamma_{p}$ being the linearized version) on the $F$-isocrystal at $x$ arising from the crystalline local system $\rho_{p}$; Kisin’s integral $p$-adic Hodge theory ([12, §2]) gives the precise well-definedness property of $\delta$. The key question remaining seems to be (for $x$ not in the $\mu$-ordinary locus, where there is no issue) finding the actual rational representative $\gamma\in G(\mathbb{Q})$; we hope to return to this problem (and the verification of the remaining conditions) in future work.
@@ 2. Recollection on companions
For a smooth variety $X$ over a finite field $k\supset\mathbb{F}_{p}$ and a geometric point $\xi$ of $X$, set $\pi_{1}(X)=\pi_{1}(X,\xi)$. Drinfeld proves ([7, Theorem 1.4.1]) that for any semisimple, not necessarily connected, group $H$ over $\overline{\mathbb{Q}}$, any prime $\lambda$ not above $p$ of $\overline{\mathbb{Q}}$, and any continuous representation $\rho_{\lambda}\colon\pi_{1}(X)\to H(\overline{\mathbb{Q}}_{\lambda})$ with Zariski-dense image, $\rho_{\lambda}$ has for any other place $\lambda^{\prime}\nmid p$ a companion
@@
also with Zariski-dense image. This $\rho_{\lambda\leadsto\lambda^{\prime}}$ has the weak companion property that for all closed points $x\in X$, $\rho_{\lambda}(\mathrm{Frob}_{x})$ and $\rho_{\lambda\leadsto\lambda^{\prime}}(\mathrm{Frob}_{x})$ define a common element of the GIT quotient (space of conjugacy classes) $[H{\,\!\sslash\!\,}H](\overline{\mathbb{Q}})$ (inside the ¯$\overline{\mathbb{Q}}_{\lambda}$ and ¯$\overline{\mathbb{Q}}_{\lambda^{\prime}}$ points, respectively). It also has a deeper strong companion property that characterizes it uniquely even in the absence of the implication “local Frobenius conjugacy implies global conjugacy.” Namely, for fixed $\lambda$, let $\hat{\Pi}_{\lambda}$ be the inverse limit over the targets of all Zariski-dense continuous homomorphisms $\pi_{1}(X,\xi)\to H(\overline{\mathbb{Q}}_{\lambda})$. Let $\widetilde{X}$ be a “universal cover” of $X$, i.e. a pro-object of finite étale covers that pro-represents a fiber functor on $\mathrm{F\acute{E}t}(X)$, let $\Pi=\operatorname{Aut}(\widetilde{X}/X)$, and choose a geometric point $\tilde{\xi}$ of $\widetilde{X}$ over our geometric point $\xi$. This choice induces an isomorphism $\Pi\cong\pi_{1}(X,\xi)^{\mathrm{op}}$, or (composing with $g\mapsto g^{-1}$) if one prefers an isomorphism $\Pi\cong\pi_{1}(X,\xi)$. $\Pi$ contains a canonical subset $\Pi_{\mathrm{Frob}}$ of Frobenius elements $\{\mathrm{Frob}_{\tilde{x}}^{n}:\tilde{x}\in|\widetilde{X}|,n\geq 1\}$ coming from the classical theory of decomposition groups, and our chosen isomorphism transfers this subset to $\pi_{1}(X,\xi)$. The groups $\hat{\Pi}_{\lambda}$ descend, uniquely up to inner automorphisms by the identity component, to groups $\hat{\Pi}_{(\lambda)}$ over $\overline{\mathbb{Q}}$. Drinfeld shows that for two places $\lambda$, $\lambda^{\prime}$, there is a unique isomorphism $\hat{\Pi}_{(\lambda)}\cong\hat{\Pi}_{(\lambda^{\prime})}$, modulo inner automorphisms by the identity component, preserving the canonical diagrams
$\Pi_{\mathrm{Frob}}\to[\hat{\Pi}_{(\lambda)}{\,\!\sslash\!\,}\hat{\Pi}^{0}_{(\lambda)}](\overline{\mathbb{Q}})\to\Pi=\pi_{0}(\hat{\Pi}_{(\lambda)}).$
@@
This isomorphism produces the companions $\rho_{\lambda\to\lambda^{\prime}}$. We will not make explicit use of this stronger companion property, but we will note when (as will be the case) our assertions of compatibility have this stronger meaning.
@@
In the present paper we will need an essentially equivalent variant with reductive monodromy groups. Drinfeld presents one such in [7, §6], where he defines a group $\Pi^{\mathrm{mot}}$ over $\overline{\mathbb{Q}}$ that similarly allows for the construction of weak or strong companions for continuous homomorphisms $\rho_{\lambda}\colon\pi_{(}X)\to H(\overline{\mathbb{Q}}_{\lambda})$ with reductive (not necessarily connected) algebraic monodromy group such that in each representation of $H$ and for all $x\in|X|$, the eigenvalues of $\mathrm{Frob}_{x}$ are all #$\#\kappa(x)$-Weil numbers. We explain another variant here.
@@
Let $\rho_{\lambda}\colon\pi_{1}(X)\to H(\overline{\mathbb{Q}}_{\lambda})$ be a continuous homomorphism with Zariski-dense image in a reductive (not necessarily connected) group $H/\overline{\mathbb{Q}}$. Assume that $\rho_{\lambda}$ is plain of characteristic $p$ in the following sense (c.f. [4, §4]): for all (finite-dimensional algebraic) representations $r$ of $H$ and all $x\in|X|$, the eigenvalues of $r(\rho_{\lambda}(\mathrm{Frob}_{x}))$ lie in ¯¯$\overline{\mathbb{Q}}\subset\overline{\mathbb{Q}}_{\lambda}$ and are $v$-adic units for all finite places $v$ not dividing $p$. Then $\rho_{\lambda}$ has a strong $\lambda^{\prime}$-companion for any finite place $\lambda^{\prime}$ of $\overline{\mathbb{Q}}$.
@@
The proof will explain the meaning of the final conclusion. Following [7, §6], let $\hat{\Pi}_{\lambda}^{\mathrm{red}}$ be the pro-reductive $\lambda$-adic completion of $\Pi$. By Proposition 3.5.2 of loc. cit., the canonical map
$\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}\times_{\widehat{\mathbb{Z}}}\mathrm{Hom}(\overline{\mathbb{Z}}_{\lambda}^{\times},\mathbb{G}_{m})$
is an isomorphism. (The two maps are the canonical projection $\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}$ and the map induced by
¯$\overline{\mathbb{Z}}_{\lambda}^{\times}\xrightarrow{\sim}\mathrm{Hom}_{\mathrm{cts}}(\widehat{\mathbb{Z}},\overline{\mathbb{Q}}_{\lambda}^{\times})\to\mathrm{Hom}_{\mathrm{cts}}(\pi_{1}(X),\overline{\mathbb{Q}}_{\lambda}^{\times})=\mathrm{Hom}(\hat{\Pi}_{\lambda}^{\mathrm{red}},\mathbb{G}_{m}),$
the middle arrow being pullback along $\pi_{1}(X)\to\mathrm{Gal}(\bar{k}/\mathbb{F}_{p})\simeq\widehat{\mathbb{Z}}$.) Replacing ¯$\overline{\mathbb{Z}}_{\lambda}^{\times}$ by its subgroup $\mathcal{W}_{p}\subset\overline{\mathbb{Q}}^{\times}\cap\overline{\mathbb{Z}}_{\lambda}^{\times}$ (intersection inside ¯$\overline{\mathbb{Q}}_{\lambda}^{\times}$) of all $p$-Weil numbers, Drinfeld obtains a pro-reductive group scheme quotient $\hat{\Pi}_{\lambda}^{\mathrm{red}}\twoheadrightarrow\hat{\Pi}_{\lambda}^{\mathrm{mot}}$ whose finite-dimensional representations are those of $\hat{\Pi}_{\lambda}^{\mathrm{red}}$ such that the corresponding $\lambda$-adic representation has, for all $x\in|X|$, $\mathrm{Frob}_{x}$-eigenvalues that are $q_{x}$-Weil numbers. Exactly the same construction applies if instead of $\mathcal{W}_{p}$ we take the subgroup

### v2

@@ 1. Introduction
    For all closed points $x\in\mathcal{S}_{K_{0},s}[1/\ell]$, the class of $\rho_{\ell}(\mathrm{Frob}_{x})$ in $[G{\,\!\sslash\!\,}G](\overline{\mathbb{Q}}_{\ell})$ lies in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$ and is independent of $\ell$ (not equal to the residue characteristic of $x$). [^1] In fact, $\rho_{\ell}(\mathrm{Frob}_{x})$ is semisimple, by [4, Theorem 7.1]
@@
Theorem 1.1 results from combining our previous work [17] with the work of Bakker-Shankar-Tsimerman ([4]) showing that all Shimura varieties have canonical integral models away from a finite set of bad primes (essentially the $N$ of Theorem 1.1). We should clarify the logical relation between these three papers: [4] makes use of Part (1) of Theorem 1.1 (see Lemma 5.1, Theorem 5.2 of loc. cit.), which we established for $\rho_{\ell}^{\mathrm{ad}}$ in [17, Theorem 2.3]. In §3, we give a proof of Part (1) of Theorem 1.1 in general that does not depend on the two earlier papers. Using their main result on integral models along with our main theorem from [17] (the analogue of Part (2) of Theorem 1.1 for $\rho_{\ell}^{\mathrm{ad}}$), Bakker-Shankar Tsimerman prove ([4, Theorem 9.7]) that for $v\nmid N$, every $\mu$-ordinary point of the special fiber $\mathcal{S}_{v}$ lifts to a characteristic zero special point. [^2] To be precise, [4] use our earlier result for $\rho_{\ell}^{\mathrm{ad}}$ to cover the non-abelian type (or the real rank at most 1) cases; in abelian type, [14, Corollary 2.3.1] establishes the desired compatibility. We in turn use this result on CM lifting of $\mu$-ordinary points to establish Parts (2) and (3) of Theorem 1.1.
@@
We now summarize each section and the arguments therein. In §2, we recall results of Drinfeld ([8]) on Deligne’s companions conjecture; we need his work on companions for representations valued in not necessarily connected reductive (not only semisimple) groups, and so we elaborate slightly on the results in [8, §6]; we also include a discussion of overconvergent $F$-isocrystal companions. In §3, we show (Corollary 3.4) that Part (1) of Theorem 1.1 holds: there is an integer $N$ such that for all $\ell$, all $\rho_{\ell}$ are unramified away from $N\ell$. For the $G^{\mathrm{ad}}$-representations, we showed this in [17, Proposition 2.3] by an abstract argument using superrigidity and Deligne’s conjecture. When $G$ is not an adjoint group, descents to $\pi_{1}(S_{K_{0},s})$ of a given representation of $\pi_{1}^{\mathrm{top}}(S_{K_{0},s}(\mathbb{C}))$ (extended to $\pi_{1}((S_{K_{0},s})_{\overline{\mathbb{Q}}})$ using integrality) are not unique, and it is easy to choose descents for varying $\ell$ that have no common finite bad set $N$ as in Part (1) of the Theorem. We need to use the fact that our descents arise from the family of canonical models; it turns out that all we need in addition to the arguments in the adjoint case is that there is a number field point $y$ of $S_{K_{0},s}$ for which the specialized local systems $\rho_{\ell}|_{y}$ are unramified outside $N_{y}\ell$ for some integer $N_{y}$ independent of $\ell$. In §4 we prove Parts (2) and (3) of Theorem 1.1. Given the result of §3, from [4, Theorem 9.2] we deduce that on the $\mu$-ordinary locus of the special fiber $\mathcal{S}_{v}$ ($v\nmid N\ell\ell^{\prime}$), $\rho_{\ell}$ and $\rho_{\ell^{\prime}}$ satisfy the companions property (2). Using the companions construction (§2) and Zariski-density of the $\mu$-ordinary locus, we can deduce the companions property on all of $\mathcal{S}_{v}$ via a strong multiplicity 1 property for Galois representations. The crystalline compatibility works analogously but requires as additional input the overconvergent ($G$-valued) F-isocrystals associated to the crystalline (by Esnault-Groechenig: [20, Appendix]) local systems $\rho_{p}$ in our joint work [10].
@@
We end this introduction by remarking on what we do not do, and what we hope to pursue in future work. We view our results in the framework of the Langlands-Rapoport conjecture, as a significant step toward associating Kottwitz triples to mod $p$ points on Shimura varieties; in abelian type, Kisin has proven the Langlands-Rapoport conjecture ([14]), but this remains a wide-open and fundamental problem in non-abelian type. We further note that Kottwitz triples are not only a useful intermediary in work on the full Langlands-Rapoport conjecture (as in loc. cit.) but are in principle a sufficient intermediary to calculate the zeta function of the Shimura variety; indeed, this is how Kottwitz first employed them in [15]. Recall ([14, §4.3]) that a Kottwitz triple over $k=\mathbb{F}_{p^{r}}$ is a triple $(\gamma_{0},(\gamma_{\ell})_{\ell\neq p},\delta)$ where
@@
  • After possibly replacing $k$ by a finite extension, there is an inner twisting $I$ of $I_{0}:=\mathrm{Cent}_{G}(\gamma_{0})$ such that $I\otimes_{\mathbb{Q}}\mathbb{R}$ is anisotropic mod center, and for all primes $\ell$, $I\otimes_{\mathbb{Q}}\mathbb{Q}_{\ell}$ is isomorphic to $I_{\ell}$ as an inner twist of $I_{0}\otimes_{\mathbb{Q}}\mathbb{Q}_{\ell}$ (here $I_{\ell}$ for $\ell\neq p$ is $\mathrm{Cent}_{G_{\mathbb{Q}_{\ell}}}(\gamma_{\ell})$, and for $\ell=p$ it is the Frobenius-centralizer: see [14, §4.3.1] for details).
To every $k$-point $x\in\mathcal{S}(k)$ ($p\nmid N$) we associate a triple $(\gamma_{0},(\gamma_{\ell})_{\ell\neq p},\delta)$, except that we have only constructed $\gamma_{0}$ in general as an element of $[G{\,\!\sslash\!\,}G](\mathbb{Q})$. With that caveat, we have verified the first and second desiderata for the triple: $\gamma_{\ell}$ is of course $\rho_{\ell}(\mathrm{Frob}_{x})$, and $\delta$ arises from the absolute Frobenius ($\gamma_{p}$ being the linearized version) on the $F$-isocrystal at $x$ arising from the crystalline local system $\rho_{p}$; Kisin’s integral $p$-adic Hodge theory ([13, §2]) gives the precise well-definedness property of $\delta$. The key question remaining seems to be (for $x$ not in the $\mu$-ordinary locus, where there is no issue) finding the actual rational representative $\gamma\in G(\mathbb{Q})$; we hope to return to this problem (and the verification of the remaining conditions) in future work.
@@ 2. Recollection on companions
For a smooth connected variety $X$ over $\mathbb{F}_{p}$ and a geometric point $\xi$ of $X$, set $\pi_{1}(X)=\pi_{1}(X,\xi)$. Drinfeld proves ([8, Theorem 1.4.1]) that for any semisimple, not necessarily connected, group $H$ over $\overline{\mathbb{Q}}$, any prime $\lambda$ not above $p$ of $\overline{\mathbb{Q}}$, and any continuous representation $\rho_{\lambda}\colon\pi_{1}(X)\to H(\overline{\mathbb{Q}}_{\lambda})$ with Zariski-dense image, $\rho_{\lambda}$ has for any other place $\lambda^{\prime}\nmid p$ a companion
@@
also with Zariski-dense image. This $\rho_{\lambda\leadsto\lambda^{\prime}}$ has the weak companion property that for all closed points $x\in X$, $\rho_{\lambda}(\mathrm{Frob}_{x})$ and $\rho_{\lambda\leadsto\lambda^{\prime}}(\mathrm{Frob}_{x})$ define a common element of the GIT quotient (space of conjugacy classes) $[H{\,\!\sslash\!\,}H](\overline{\mathbb{Q}})$ (inside the ¯$\overline{\mathbb{Q}}_{\lambda}$ and ¯$\overline{\mathbb{Q}}_{\lambda^{\prime}}$ points, respectively); here and throughout we write $\mathrm{Frob}_{x}$ for a geometric Frobenius element associated to the residue field $\kappa(x)$. It also has a deeper strong companion property that characterizes it uniquely even in the absence of the implication “local Frobenius conjugacy implies global conjugacy.” We give Drinfeld’s formulation in terms of a “universal cover” $\widetilde{X}$ of $X$. Let $\widetilde{X}$ be a pro-object of the category of finite étale covers of $X$ that pro-represents a fiber functor on $\mathrm{F\acute{E}t}(X)$, let $\Pi=\operatorname{Aut}(\widetilde{X}/X)$, and choose a geometric point $\tilde{\xi}$ of $\widetilde{X}$ over our geometric point $\xi$. This choice induces an isomorphism $\Pi\cong\pi_{1}(X,\xi)^{\mathrm{op}}$ and (composing with $g\mapsto g^{-1}$) an isomorphism $\Pi\cong\pi_{1}(X,\xi)$, which will allow the assertions on existence of companions to be transferred to $\pi_{1}(X)$.
Now, for fixed $\lambda$, let $\hat{\Pi}_{\lambda}$ be the inverse limit over the targets of all Zariski-dense continuous homomorphisms $\Pi\to H(\overline{\mathbb{Q}}_{\lambda})$ to semisimple groups $H$. $\Pi$ contains a canonical subset of Frobenius elements $\Pi_{\mathrm{Frob}}=\{\mathrm{Frob}_{\tilde{x}}^{n}:\tilde{x}\in|\widetilde{X}|,n\geq 1\}$ coming from the classical theory of decomposition groups, and our chosen isomorphism transfers this subset to $\pi_{1}(X)$ as the union of conjugacy classes of positive powers of geometric Frobenii at all closed points. The groups $\hat{\Pi}_{\lambda}$ descend, uniquely up to inner automorphisms by the identity component, to groups $\hat{\Pi}_{(\lambda)}$ over $\overline{\mathbb{Q}}$. Drinfeld shows that for two places $\lambda$, $\lambda^{\prime}$, there is a unique isomorphism $\hat{\Pi}_{(\lambda)}\cong\hat{\Pi}_{(\lambda^{\prime})}$, modulo inner automorphisms by the identity component, preserving the canonical diagrams
$\Pi_{\mathrm{Frob}}\to[\hat{\Pi}_{(\lambda)}{\,\!\sslash\!\,}\hat{\Pi}^{0}_{(\lambda)}](\overline{\mathbb{Q}})\to\Pi=\pi_{0}(\hat{\Pi}_{(\lambda)}).$ (1)
@@
This isomorphism produces the companions $\rho_{\lambda\leadsto\lambda^{\prime}}$: a homomorphism $\hat{\Pi}_{\lambda}\to H_{\overline{\mathbb{Q}}_{\lambda}}$ induces a homomorphism $\hat{\Pi}_{(\lambda)}\to H$ (unique up to $\hat{\Pi}_{(\lambda)}^{0}$-conjugation), which can then be transported to its $\lambda^{\prime}$-adic version. We will not make explicit use of this stronger companion property, but we will note when (as will be the case) our assertions of compatibility have this stronger meaning.
@@
In the present paper we will need a variant with reductive monodromy groups; we will also need the companion construction (in the reductive case) when $\lambda^{\prime}|p$. In the case $\lambda,\lambda^{\prime}\nmid p$, Drinfeld presents one reductive variant in [8, §6], where he defines a group $\Pi^{\mathrm{mot}}$ over $\overline{\mathbb{Q}}$ that similarly allows for the construction of weak or strong companions for continuous homomorphisms $\rho_{\lambda}\colon\Pi\to H(\overline{\mathbb{Q}}_{\lambda})$ with reductive (not necessarily connected) algebraic monodromy group such that in each representation of $H$ and for all $x\in|X|$, the eigenvalues of $\mathrm{Frob}_{x}$ are all #$\#\kappa(x)$-Weil numbers. We will explain this along with a version that produces companions for local systems that are plain of characteristic $p$ (rather than “motivic,” i.e. a direct sum of pure) in the sense of [5].
Following [8, §6], for $\lambda\nmid p$ let $\hat{\Pi}_{\lambda}^{\mathrm{red}}$ be the pro-reductive $\lambda$-adic completion of $\Pi$. Now let $\lambda|p$, and let $\operatorname{F-Isoc}^{\dagger}(X)$ be the category of overconvergent $F$-isocrystals on $X$. Its ¯$\overline{\mathbb{Q}}_{\lambda}$-linearization $\operatorname{F-Isoc}^{\dagger}\otimes_{\mathbb{Q}_{p}}\overline{\mathbb{Q}}_{\lambda}$ is a neutral Tannakian category over ¯$\overline{\mathbb{Q}}_{\lambda}$, and choosing a fiber functor (for instance, by pulling back to a closed point: see [10, Example 3.1 ff.]) we obtain a Tannakian group $\pi_{1}^{\operatorname{F-Isoc}^{\dagger}}(X)$. As in the $\lambda\nmid p$ case, we write $\hat{\Pi}_{\lambda}$ and $\hat{\Pi}_{\lambda}^{\mathrm{red}}$ for the maximal pro-semisimple and maximal pro-reductive quotients of $\pi_{1}^{\operatorname{F-Isoc}^{\dagger}}(X)$.
We have the following maps:
  • Suppose $\lambda\nmid p$. Then $\Xi_{\lambda}\colon\overline{\mathbb{Z}}_{\lambda}^{\times}\to\mathrm{Hom}(\hat{\Pi}_{\lambda}^{\mathrm{red}},\mathbb{G}_{m})$ is constructed as the composition
    $\Xi_{\lambda}\colon\overline{\mathbb{Z}}_{\lambda}^{\times}\simeq\mathrm{Hom}_{\mathrm{cts}}(\hat{\mathbb{Z}},\overline{\mathbb{Q}}_{\lambda}^{\times})\to\mathrm{Hom}_{\mathrm{cts}}(\Pi,\overline{\mathbb{Q}}_{\lambda}^{\times})=\mathrm{Hom}(\hat{\Pi}_{\lambda}^{\mathrm{red}},\mathbb{G}_{m}),$
    where the second map is precomposition with the canonical homomorphism $\Pi\to\hat{\mathbb{Z}}$ (under which $\mathrm{Frob}_{\tilde{x}}$ maps to $[\kappa(x):\mathbb{F}_{p}]$).
  • Suppose $\lambda|p$. Then $\Xi_{\lambda}\colon\overline{\mathbb{Q}}_{\lambda}^{\times}\to\mathrm{Hom}(\hat{\Pi}_{\lambda}^{\mathrm{red}},\mathbb{G}_{m})$ is given by interpreting an element of ¯$\overline{\mathbb{Q}}_{\lambda}^{\times}$ as a rank-1 object of $\operatorname{F-Isoc}(\mathrm{Spec}(\mathbb{F}_{p}))\otimes_{\mathbb{Q}_{p}}\overline{\mathbb{Q}}_{\lambda}$ (see [10, Example 3.1]) and then pulling back along $X\to\mathrm{Spec}(\mathbb{F}_{p})$.
These homomorphism (dualized) and the canonical projections $\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}$ induce the following isomorphisms:
  • Suppose $\lambda\nmid p$. Then
    $\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\overline{\mathbb{Z}}_{\lambda}^{\times},\mathbb{G}_{m})$
    is an isomorphism by [8, Proposition 3.5.2]
  • Suppose $\lambda|p$. Then
    $\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\overline{\mathbb{Q}}_{\lambda}^{\times},\mathbb{G}_{m})$
    is an isomorphism by Lemma 2.1 below.
I owe the following lemma to Jake Huryn’s write-up in a draft version of [10]:
@@
Suppose $\lambda|p$. Then the map
$\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\overline{\mathbb{Q}}_{\lambda}^{\times},\mathbb{G}_{m})$
just described is an isomorphism.
@@
We use the criterion of [8, Lemma 3.5.1]. ¯$\overline{\mathbb{Q}}_{\lambda}^{\times}/\mu_{\infty}$ is a $\mathbb{Q}$-vector space, so it suffices to show that for all finite index subgroups $G^{\prime}$ of $\hat{\Pi}_{\lambda}^{\mathrm{red}}$, the map
¯$\overline{\mathbb{Q}}_{\lambda}^{\times}/\mu_{\infty}\to\mathrm{Hom}(G^{\prime},\mathbb{G}_{m})/\mathrm{torsion}$
induced by $\Xi_{\lambda}$ is an isomorphism. First, by [7, Proposition B.7.6] any finite index subgroup $G^{\prime}\subset\hat{\Pi}_{\lambda}^{\mathrm{red}}$ is the image of the injection $\hat{\Pi}_{\lambda,X^{\prime}}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}^{\mathrm{red}}$, where $X^{\prime}=\widetilde{X}/U\to X$ is determined by taking the image $U$ of $G^{\prime}$ in $\Pi$ (and $\hat{\Pi}_{\lambda,X^{\prime}}^{\mathrm{red}}$ denotes the analogous pro-reductive completion of $\pi_{1}^{\operatorname{F-Isoc}^{\dagger}}(X^{\prime})$). If $\Xi_{\lambda}(\alpha)|_{X^{\prime}}$ is torsion, then the $F$-isocrystal on $\mathrm{Spec}(\mathbb{F}_{p})$ associated to $\alpha$ is already torsion after passing to some finite field extension $k/\mathbb{F}_{p}$ (e.g., the field of constants of $X^{\prime}$), where the linearized Frobenius acts as $\alpha^{[k:\mathbb{F}_{p}]}$. Thus $\alpha\in\mu_{\infty}$ is torsion.
For surjectivity, we use the result of Crew ($\dim(X)=1$) and Abe (the general case) that up to a twist pulled back from the ground field any rank 1 overconvergent $F$-isocrystal on a finite-type $\mathbb{F}_{p}$-scheme is torsion ([2, Lemma 6.1]). Let $\mathcal{E}$ be the rank 1 overconvergent $F$-isocrystal on $X^{\prime}$ corresponding to an element of $\mathrm{Hom}(\hat{\Pi}_{\lambda,X^{\prime}}^{\mathrm{red}},\mathbb{G}_{m})$; then there exists a rank 1 $F$-isocrystal $\mathcal{F}$ on $X^{\prime}$ that is pulled back from $\mathrm{Spec}(\mathbb{F}_{p})$ such that $\mathcal{E}\otimes\mathcal{F}$ is torsion. Thus, up to twisting by torsion objects $\mathcal{E}$ lies in the image of $\Xi_{\lambda}$, as desired.
Replacing ¯$\overline{\mathbb{Z}}_{\lambda}^{\times}$ by its subgroup $\mathcal{W}_{p}\subset\overline{\mathbb{Q}}^{\times}\cap\overline{\mathbb{Z}}_{\lambda}^{\times}$ (intersection inside ¯$\overline{\mathbb{Q}}_{\lambda}^{\times}$) of all $p$-Weil numbers, Drinfeld obtains a pro-reductive group scheme quotient $\hat{\Pi}_{\lambda}^{\mathrm{red}}\twoheadrightarrow\hat{\Pi}_{\lambda}^{\mathrm{mot}}$ whose finite-dimensional representations are those of $\hat{\Pi}_{\lambda}^{\mathrm{red}}$ such that the corresponding $\lambda$-adic representation has, for all $x\in|X|$, $\mathrm{Frob}_{x}$-eigenvalues that are #$\#\kappa(x)$-Weil numbers. Exactly the same construction applies if instead of $\mathcal{W}_{p}$ we take the subgroup